Circular No. 8976

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1309 SCORPII R. J. Rudy, D. K. Lynch, R. W. Russell, and B. Kaneshiro, The

Aerospace Corporation; and M. Sitko and H. Hammel, Space Science Institute, report 0.8- to 2.4-μm spectroscopy of V1309 Sco (N Sco 2008; cf. IAUC 8972) on Sept. 5 UT using the Spex instrument at the Infrared Telescope Facility. V1309 Sco is at a very early stage of development, mimicking an F-giant spectrum and showing weak emission lines for only a few of the lower series Pashchen transitions and some of the stronger transitions of Fe II. The spectrum is dominated by aborption features including most of the higher Pashcen series, all of the Brackett lines, the Ca II triplet, and lines of C I and N I. The presence of iron emission at this early stage indicates that V1309 Sco is an “Fe II”-type nova. This, plus the narrowness of the lines (full-width-at-half-maximum ∼ 300 km/s) and the predominance of the absorption spectrum suggest that the V1309 Sco eruption resulted from the ejection of a massive, low-velocity shell by a carbon-oxygen white dwarf.

Additional magnitude estimates of V1309 Sco (visual unless otherwise noted): Aug. 26.52 UT, [11.7 (N. J. Brown, Ellenbrook, Western Australia, 200-mm f/2.8 telephoto lens + T-Max 400 film); 30.019, [12.4 (W. Liller, Viña del Mar, Chile, Technical Pan film + orange filter); Sept. 3.952, 9.3 (J. G. de S. Aguiar, Campinas, Brazil); 4.955, 8.5 (Aguiar); 5.940, 8.1 (Aguiar); 6.015, 7.5 (Liller, as above).

(136108) HAUMEA The IAU’s Committee on Small Body Nomenclature and Working

Group for Planetary System Nomenclature have agreed to give to the transneptunian dwarf planet, or ‘plutoid’, (136108) 2003 EL_{61} (cf. IAUC 8577) the name “Haumea”, with the citation given on MPC 63878. At the same time, the IAU is giving the name Hi’iaka to (136108) Haumea I — the larger, outer satellite (cf. IAUC 8577) — and the name “Namaka” to (136108) Haumea II — the smaller, inner satellite (cf. IAUC 8636).

COMETS 201P, 202P, 203P, 204P, 205P The following permanent numbers have been assigned (cf. MPC 63768)

to newly recovered comets: 201P/2008 Q4 (LONEOS; cf. IAUC 8970); 202P/2008 R2 (Scotti; cf. IAUC 8971); 203P/2008 R4 (Korlević; cf. IAUC 8974); 204P/2008 R5 (LINEAR-NEAT; cf. IAUC 8974); 205P/2008 R6 (Giacobini; cf. IAUC 8975).

2008 September 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green